OpenStudy (anonymous):

how to find the distances between lines? let's say 2x+3y=6 and 4x+6y=8 ?

OpenStudy (paxpolaris):

first of all the two lines have to be parallel (which these are) write them in point slope form...

OpenStudy (anonymous):

If the lines are parallel, you need to measure a perpendicular distance between them.

OpenStudy (anonymous):

okay and what if lines are not parallel?

OpenStudy (anonymous):

If they aren't parallel and in fact cross somewhere, then "distance between them" depends on where you want to check... could be big, could be nothing at the intersection.

OpenStudy (anonymous):

@jake how to find perpendicular distance ?

OpenStudy (anonymous):

Find the slope of each to see if they are parallel (same slope)

OpenStudy (anonymous):

that was just an example...okay let's say 2x+4y=9 and 2x+4y=10 now?

OpenStudy (anonymous):

Sorry, internet lag here... my responses are showing up with a lot of delay.

OpenStudy (anonymous):

that's okay

OpenStudy (amistre64):



OpenStudy (anonymous):

@amistre64 has the diagram (darn slow internet!!)

OpenStudy (anonymous):

@amistre64 do i need to plot that again and again...?

OpenStudy (anonymous):

Do you really need to go with "sin(a)" to get this though? Agree that works, but is there a non-trig way?

OpenStudy (anonymous):

is there any formula or something

OpenStudy (amistre64):

its a general setup to notice what parts you can use

OpenStudy (amistre64):

the slope of a line is the tangent of the angle between x axis, so "a" in this case is 90-tan^(-1) (slope) and b is the difference in y intercepts

OpenStudy (amistre64):

which should be the same as the difference between the constants used

OpenStudy (anonymous):

@erica.d if you are comfortable with that approach, that works well. Otherwise, you may have some extra work to find points on the "connector line" where it intersects your two original lines, then you could do distance formula with those points.

OpenStudy (anonymous):

@JakeV8 thanks...i was just wondering if i could use some formula

OpenStudy (anonymous):

Would be nice, wouldn't it ? :)

OpenStudy (amistre64):

using: 2x+4y=9 2x+4y=10 ; 10-9 = 1 m = -2/4 = -.5 1*sin(90-arctan(-.5)) is what im thinking :)

OpenStudy (anonymous):

Do you know the distance formula between two points?

OpenStudy (amistre64):

might have to reconsider my "constants" 10/4 - 9/4 = 1/4 ... not 1 sooo

OpenStudy (anonymous):

@amistre64 I'm taking notes :) I hadn't thought of your approach :)

OpenStudy (anonymous):

thanks @amistre64 @JakeV8

OpenStudy (amistre64):

:) i get .15 if we have something to compare with

OpenStudy (amistre64):

im thinking i can make it simpler with a cosine .... b cos(arctan(m))

OpenStudy (amistre64):



hero (hero):

Yes, there is a formula for this

hero (hero):

You can read more about it here: http://tinyurl.com/9lk9cwu

OpenStudy (anonymous):

@Hero That's a great page... thanks for the link!

hero (hero):

@JakeV8, that's my blog. I wrote it.

OpenStudy (anonymous):

Writing it is much more awesome than "googling" it... you can find help lots of places, but to create your own is pretty cool.